OpenStudy (anonymous):

A company had sales of $500,000 in 1996 and sales of $720,000 in 1998. Use the midpoint formula to find the company's sales in 1997.

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

((2+4)/2,(3+1)/2) = (3,2)

OpenStudy (anonymous):

you need a point, so it will be in the form (x,y). You need to find the midpoint for the x values first (2+4) / 2 = x then for the y values (3+1) / 2 = y and then your answer is (x,y)

OpenStudy (anonymous):

K?

OpenStudy (anonymous):

how though? please explain it in more detail.

OpenStudy (anonymous):

Medal

OpenStudy (anonymous):

First tell me the answer and explain it first then you get your medal :)

OpenStudy (mathstudent55):

Use (1996, $500,000) as one point, and (1998, $720,000) as the other point. To find the midpoint of two points, find the average of the x-values and the average of the y-values. For example, what is the midpoint of (1, 2) and (5, 6)? ( (1 + 5)/2, (2 + 6)/2 ) = (3, 4)

OpenStudy (anonymous):

x1=1997,x2=1998 y1=500,000,y2=720000 midpoint =(1996+1998/2 , 500,000+720000/2)=(1997,610000) midpoint=(1997,610000)